4x-y=4
5x+5y=9
Solve using the substitution method

Answer by jorel1380(3719)

4x-y=4
4x-4=y
5x+5y=9
5x+5(4x-4)=9
5x+20x-20=9
25x=29
x=29/25
y=16/25